OpenStudy (anonymous):

Integration of cotx(cosecx-1)e^x dx plz help am not able to solve it :'(

OpenStudy (kc_kennylau):

This should not be in the school :P

OpenStudy (anonymous):

it is :P cbse question

OpenStudy (kc_kennylau):

If WolframAlpha says that it uses new functions, it is not in school syllabus :P

OpenStudy (anonymous):

@kc_kennylau -even my teacher was not able to solve it it's just one mark question but still no one is able to get the answr if that was not a cbse sample paper question then 100 % it was incorrect question

OpenStudy (anonymous):

lol it is

OpenStudy (anonymous):

again this question has been posted in yahoo answer so i was thinking this time may be i will get the answer but still i m not able to solve it lol there must be something in thsi question that's y people are not able to sovle it

OpenStudy (kc_kennylau):

yes, it was incorrect question, it uses newly defined functions

OpenStudy (anonymous):

yeah correct version shoudl be like tjhis e^x(cotx-cosec^2x)dx

OpenStudy (kainui):

So, do you recognize any kind of relationship between cot(x) and -csc^2(x)?

OpenStudy (kainui):

Similarly, d/dx(tanx)=sec^2(x). This is fairly similar, but with cotangent and cosecant.

OpenStudy (anonymous):

Using the correct versions its a simple integration by parts problem I believe.

OpenStudy (kainui):

@malevolence19 it's much easier than that. It's a simple substitution problem.

OpenStudy (anonymous):

But you have: \[\int e^x (\cot x-\csc^2 x)dx\] With the e^x you still have integration by parts yes?

OpenStudy (anonymous):

Finding the integrals of cotangent and cosecant squared are easy enough but I'm not seeing what substitution would help honestly, maybe I'm tired. :/

OpenStudy (kainui):

So if you you pick:\[u=e^x \cot x\]\[du=(e^x \cot x - e^x \csc^2x) dx\]